OpenStudy (anonymous):

HELP PLEASE! What do you know about the value of b in the function y=ax^2+bx+c when the x-coordinate of the vertex is an integer?

OpenStudy (mathstudent55):

The x-coordinate of the vertex is \(x = \dfrac{-b}{2a} = -\dfrac{1}{2} \times \dfrac{b}{a}\)

OpenStudy (amistre64):

let n be an integer value n = -b/(2a) solve for b, what becomes evident about b? what type of number is it?